# Determinantal Point Process Attention Over Grid Cell Code Supports Out of Distribution Generalization

### 18 Jan 2024 v2

@@ 4.2 Comparison models
To evaluate how the grid cell code coupled with DPP-A compares with other architectures and approaches to generalization, and the extent to which each of these components contributed to performance of the model, we compared it with several alternative models for performing the analogy and arithmetic tasks. First we compared it with the temporal context normalization (TCN) model (Webb et al., 2020) (see Section 3), but modified so as to use the grid cell code as input. We passed the grid cell embeddings for each input through a shared feedforward encoder which consisted of two fully connected layers with 256 units per layer. ReLU nonlinearities were used in both the layers. The final embedding was generated with a linear layer of 256 units. TCN was applied to these embeddings and then passed as a sequence for each candidate completion to the inference module. This implementation of TCN involved a learned encoder on top of the grid cell embeddings, so it is closely analogous to the original TCN.
Next, we compared our model to one that used variational dropout (Gal & Ghahramani, 2016), which is shown to be more effective in generalization compared to naive dropout (Srivastava et al., 2014). We randomly sampled a dropout mask (50% dropout), zeroing out the contribution of some of the grid cell code in the input to the inference module. We then use that locked dropout mask for every time step in the sequence. We also compared DPP-A to a model that had an additional penalty (L1 regularization) proportional to the absolute sum of the attention gates $g$ along with the task-specific loss. We experimented with different values of $\lambda$, which controlled the strength of the penalty relative to the cross entropy loss. We report accuracy values for $\lambda$ that achieved the best performance on the validation set. Accuracy values for various $\lambda$s are provided in the Appendix 7.8. Dropout and L1 regularization were chosen as a proxy for DPP-A and hence we used the same input data for fair comparison. Finally, we compared to a model that used the complete grid cell code, i.e. no DPP-A.
@@ 5.1 Analogy
We first present results on analogy task for two types of testing data, translation and scaling using two types of inference module, LSTM and transformer. We trained 3 networks for each method and report mean accuracy along with standard error of the mean. Figure 3 shows the results for the analogy task using an LSTM in the inference module. The left panel shows results for the translation regime, and the right panel shows results for the scaling regime. Both plots show accuracy on the training and validation sets, and on a series of 9 (increasingly distant) OOD generalization test regions. DPP-A (shown in blue) achieves nearly perfect accuracy on all of the test regions, considerably outperforming the other models.
@@
Figure 4: Results on analogy on each region for translation and scaling using transformer in the inference module.
@@
To test the generality of the grid cell code and DPP-A across network architectures, we also tested a transformer (Vaswani et al., 2017) in place of the LSTM in the inference module. Previous work has suggested that transformers are particularly useful for extracting structure in sequential data and has been used for OOD generalization (Saxton et al., 2019). Figure 4 shows the results for the analogy task using a transformer in the inference module. With no explicit attention (no DPP-A) over the grid cell code (show in orange), the transformer did poorly on the analogies on the test regions. This suggests that simply using a more sophisticated architecture with standard forms of attention is not sufficient to enable OOD generalization based on the grid cell code. With DPP-A (shown in blue), the OOD generalization performance of the transformer is nearly perfect for both translation and scaling. These results also demonstrate that grid cell code coupled with DPP-A can be exploited for OOD generalization effectively by different architectures.
@@ 5.2 Arithmetic
We next present results on arithmetic task for two types of operations, addition and multiplication using two types of inference module, LSTM and transformer. We trained 3 networks for each method and report mean accuracy along with standard error of the mean.
@@
Figure 6: Results on arithmetic on each region using transformer in the inference module.
@@
Figure 6 shows the results for arithmetic problems using a transformer in the inference module. With no DPP-A over the grid cell code the transformer did poorly on addition and multiplication on the test regions, achieving around 20-30% accuracy. With DPP-A, the OOD generalization performance of transformer show a pattern similar to that for LSTM: it is nearly perfect for addition and, though not as good on multiplication, nevertheless show approximately 40% better performance than the transformer multiplication.
@@ 5.3 Ablation study
Figure 7 shows the results for the analogy task (results for arithmetic task are in Appendix 7.6, Figure 11) using an LSTM in the inference module. The average accuracy on the test regions for translation and scaling using smoothed one-hots passed through an encoder (shown in green) is nearly 30% better than simple one-hot embeddings passed through an encoder (shown in orange), but both still achieve significantly lower test accuracy than grid cell code which support perfect OOD generalization.
@@
Finally, for completeness, we also carried out a set of simulations that examined the performance of networks with various embeddings (grid cell code, and one-hots or smoothed one-hots with or without a learned feedforward encoder), but no attention or regularization (i.e., neither DPP-A, transformer, nor TCN, L1 Regularization, or Dropout). Figure 8 shows the results for the different embeddings on the analogy task (results for arithmetic task are in Appendix 7.6, Figure 12). For translation (left), the average accuracy over the test regions using grid cell code (shown in blue) is nearly 25% more compared to other embeddings, which all yield performance near chance ($\sim$ 15%). For scaling (right), although other embeddings achieve higher performance than chance (except smoothed one-hots), they still achieve lower test accuracy than grid cell code. More ablation studies can be found in Appendix 7.4, 7.9, and Figure 13.
@@ 6 Discussion and future directions
We have identified how particular properties of processing observed in the brain can be used to achieve strong OOD generalization, and introduced a two-component algorithm to promote OOD generalization in deep neural networks. The first component is a structured representation of the training data, modeled closely on known properties of grid cells – a population of cells that collectively represent abstract position using a periodic code. However, despite their intrinsic structure, we find that grid cell code and standard error-driven learning alone are insufficient to promote OOD generalization, and standard approaches for preventing overfitting offer only modest gains. This is addressed by the second component, using DPP-A to implement attentional regularization over the grid cell code. DPP-A allows only a relevant and diverse subset of cells to influence downstream computation in the inference module using the statistics of the training data. For proof of concept, we started with two challenging cognitive tasks (analogy and arithmetic), and showed that the combination of grid cell code and DPP-A promotes OOD generalization across both translation and scale when incorporated into common architectures (LSTM and transformer). It is worth noting that the DPP-A attentional mechanism is different from the attentional mechanism in the transformer module, and both are needed for strong OOD generalization in the case of transformers. Use of the standard self-attention mechanism in transformers over the inputs (i.e., A, B, C and D for the analogy task) in place of DPP-A would lead to weightings of grid cell embeddings over all frequencies and phases. The objective function for the DPP-A represents an inductive bias, that selectively assigns the greatest weight to all grid cell embeddings (i.e., for all phases) of the frequency for which the determinant of the covariance matrix is greatest computed over the training space. The transformer inference module then attends over the inputs with the selected grid cell embeddings based on the DPP-A objective.
@@
The current approach has some limitations and presents interesting directions for future work. First, we derive the grid cell code from known properties of neural systems, rather than obtaining the code directly from real-world data. Here, we are encouraged by the body of work providing evidence for grid cell code in the hidden layers of neural networks in a variety of task contexts and architectures (Wei et al., 2015; Cueva & Wei, 2018; Banino et al., 2018; Whittington et al., 2020). This suggests reason for optimism that DPP-A may promote strong generalization in cases where grid cell code naturally emerge: for example, navigation tasks (Banino et al., 2018) and reasoning by transitive inference (Whittington et al., 2020). Integrating our approach with structured representations acquired from high-dimensional, naturalistic datasets remains a critical next step which would have significant potential for broader future practical applications. So too does application to more complex transformations beyond translation and scale, such as rotation, and complex forms of representations, and analogical reasoning tasks (Holyoak, 2012; Webb et al., 2023; Lu et al., 2022). Second, it is not clear how DPP-A might be implemented in a neural network. In that regard, Bozkurt et al. (2022) recently proposed a biologically plausible neural network algorithm using a weighted similarity matrix approach to implement a determinant maximization criterion, which is the core idea underlying the objective function we use for DPP-A (Equation 6), suggesting that the DPP-A mechanism we describe may also be biologically plausible. This could be tested experimentally by exposing individuals (e.g., rodents or humans) to a task that requires consistent exposure to a subregion, and evaluating the distribution of activity over the grid cells. Our model predicts that high frequency grid cells should increase their firing rate more than low frequency cells, since the high frequency grid cells maximize the determinant of the covariance matrix of the grid cell embeddings. It is also worth noting that (Frankland & Cohen, 2020) have suggested that the use of DPPs may also help explain a mutual exclusivity bias observed in human word learning and reasoning. While this is not direct evidence of biological plausibility, it is consistent with the idea that the human brain selects representations for processing that maximize the volume of the representational space, which can be achieved by maximizing the DPP-A objective function defined in Equation 6. Third, we compared grid cell code to only one-hots and place cell code. Future work could compare to a broader range of potential biological coding schemes for the overall space, e.g. boundary vector cell coding (Barry et al., 2006), band cell coding, or egocentric boundary cell coding (Hinman et al., 2019).
@@
Finally, we focus on analogies in linear spaces which limits the generality of our approach in nonlinear spaces. In that regard, there are at least two potential directions that could be pursued. One is to directly encode nonlinear structures (such as trees, rings, etc.) with grid cells, to which DPP-A could be applied as described in our model. The TEM model (Whittington et al., 2020) suggests that grid cells in the medial entorhinal may form a basis set that captures structural knowledge for such nonlinear spaces, such as social hierarchies and transitive inference when formalized as a connected graph. Another would be to use eigen-decomposition of the successor representation (Dayan, 1993), a learnable predictive representation of possible future states that has been shown by Stachenfeld et al. (2017) to provide an abstract structured representation of a space that is analogous to the grid cell code. This general purpose mechanism could be applied to represent analogies in nonlinear spaces (Frankland et al., 2019), for which there may not be a clear factorization in terms of grid cells (i.e., distinct frequencies and multiple phases within each frequency). Since the DPP-A mechanism, as we have described it, requires representations to be factored in this way it would need to be modified for such purpose. Either of these approaches, if successful, would allow our model to be extended to domains containing nonlinear forms of structure. To the extent that different coding schemes (i.e., basis sets) are needed for different forms of structure, the question of how these are identified and engaged for use in a given setting is clearly an important one, that is not addressed by the current work. We imagine that this is likey subserved by monitoring and selection mechanisms proposed to underlie the capacity for selective attention and cognitive control (Shenhav et al., 2013), though the specific computational mechanisms that underlie this function remain an important direction for future research.

### 23 Jan 2024 v3

@@ 4.2 Comparison models
To evaluate how the grid cell code coupled with DPP-A compares with other architectures and approaches to generalization, and the extent to which each of these components contributed to the performance of the model, we compared it with several alternative models for performing the analogy and arithmetic tasks. First, we compared it with the temporal context normalization (TCN) model (Webb et al., 2020) (see Section 3), but modified so as to use the grid cell code as input. We passed the grid cell embeddings for each input through a shared feedforward encoder which consisted of two fully connected layers with 256 units per layer. ReLU nonlinearities were used in both the layers. The final embedding was generated with a linear layer of 256 units. TCN was applied to these embeddings and then passed as a sequence for each candidate completion to the inference module. This implementation of TCN involved a learned encoder on top of the grid cell embeddings, so it is closely analogous to the original TCN.
Next, we compared our model to one that used variational dropout (Gal & Ghahramani, 2016), which is shown to be more effective in generalization compared to naive dropout (Srivastava et al., 2014). We randomly sampled a dropout mask (50% dropout), zeroing out the contribution of some of the grid cell code in the input to the inference module. We then use that locked dropout mask for every time step in the sequence. We also compared DPP-A to a model that had an additional penalty (L1 regularization) proportional to the absolute sum of the attention gates $g$ along with the task-specific loss. We experimented with different values of $\lambda$, which controlled the strength of the penalty relative to the cross-entropy loss. We report accuracy values for $\lambda$ that achieved the best performance on the validation set. Accuracy values for various $\lambda$s are provided in the Appendix 7.8. Dropout and L1 regularization were chosen as a proxy for DPP-A and hence we used the same input data for fair comparison. Finally, we compared to a model that used the complete grid cell code, i.e. no DPP-A.
@@ 5.1 Analogy
We first present results on the analogy task for two types of testing data, translation and scaling using two types of inference module, LSTM and transformer. We trained 3 networks for each method and report mean accuracy along with standard error of the mean. Figure 3 shows the results for the analogy task using an LSTM in the inference module. The left panel shows results for the translation regime, and the right panel shows results for the scaling regime. Both plots show accuracy on the training and validation sets, and on a series of 9 (increasingly distant) OOD generalization test regions. DPP-A (shown in blue) achieves nearly perfect accuracy on all of the test regions, considerably outperforming the other models.
@@
Figure 4: Results on analogy on each region for translation and scaling using the transformer in the inference module.
@@
To test the generality of the grid cell code and DPP-A across network architectures, we also tested a transformer (Vaswani et al., 2017) in place of the LSTM in the inference module. Previous work has suggested that transformers are particularly useful for extracting structure in sequential data and have been used for OOD generalization (Saxton et al., 2019). Figure 4 shows the results for the analogy task using a transformer in the inference module. With no explicit attention (no DPP-A) over the grid cell code (shown in orange), the transformer did poorly on the analogies on the test regions. This suggests that simply using a more sophisticated architecture with standard forms of attention is not sufficient to enable OOD generalization based on the grid cell code. With DPP-A (shown in blue), the OOD generalization performance of the transformer is nearly perfect for both translation and scaling. These results also demonstrate that grid cell code coupled with DPP-A can be exploited for OOD generalization effectively by different architectures.
@@ 5.2 Arithmetic
We next present results on the arithmetic task for two types of operations, addition and multiplication using two types of inference module, LSTM and transformer. We trained 3 networks for each method and report mean accuracy along with standard error of the mean.
@@
Figure 6: Results on arithmetic on each region using the transformer in the inference module.
@@
Figure 6 shows the results for arithmetic problems using a transformer in the inference module. With no DPP-A over the grid cell code the transformer did poorly on addition and multiplication on the test regions, achieving around 20-30% accuracy. With DPP-A, the OOD generalization performance of the transformer shows a pattern similar to that for LSTM: it is nearly perfect for addition and, though not as good on multiplication, nevertheless shows approximately 40% better performance than the transformer multiplication.
@@ 5.3 Ablation study
Figure 7 shows the results for the analogy task (results for the arithmetic task are in Appendix 7.6, Figure 11) using an LSTM in the inference module. The average accuracy on the test regions for translation and scaling using smoothed one-hots passed through an encoder (shown in green) is nearly 30% better than simple one-hot embeddings passed through an encoder (shown in orange), but both still achieve significantly lower test accuracy than grid cell code which support perfect OOD generalization.
@@
Finally, for completeness, we also carried out a set of simulations that examined the performance of networks with various embeddings (grid cell code, and one-hots or smoothed one-hots with or without a learned feedforward encoder), but no attention or regularization (i.e., neither DPP-A, transformer, nor TCN, L1 Regularization, or Dropout). Figure 8 shows the results for the different embeddings on the analogy task (results for the arithmetic task are in Appendix 7.6, Figure 12). For translation (left), the average accuracy over the test regions using grid cell code (shown in blue) is nearly 25% more compared to other embeddings, which all yield performance near chance ($\sim$ 15%). For scaling (right), although other embeddings achieve higher performance than chance (except smoothed one-hots), they still achieve lower test accuracy than grid cell code. More ablation studies can be found in Appendix 7.4, 7.9, and Figure 13.
@@ 6 Discussion and future directions
We have identified how particular properties of processing observed in the brain can be used to achieve strong OOD generalization, and introduced a two-component algorithm to promote OOD generalization in deep neural networks. The first component is a structured representation of the training data, modeled closely on known properties of grid cells – a population of cells that collectively represent abstract position using a periodic code. However, despite their intrinsic structure, we find that grid cell code and standard error-driven learning alone are insufficient to promote OOD generalization, and standard approaches for preventing overfitting offer only modest gains. This is addressed by the second component, using DPP-A to implement attentional regularization over the grid cell code. DPP-A allows only a relevant and diverse subset of cells to influence downstream computation in the inference module using the statistics of the training data. For proof of concept, we started with two challenging cognitive tasks (analogy and arithmetic), and showed that the combination of grid cell code and DPP-A promotes OOD generalization across both translation and scale when incorporated into common architectures (LSTM and transformer). It is worth noting that the DPP-A attentional mechanism is different from the attentional mechanism in the transformer module, and both are needed for strong OOD generalization in the case of transformers. Use of the standard self-attention mechanism in transformers over the inputs (i.e., A, B, C, and D for the analogy task) in place of DPP-A would lead to weightings of grid cell embeddings over all frequencies and phases. The objective function for the DPP-A represents an inductive bias, that selectively assigns the greatest weight to all grid cell embeddings (i.e., for all phases) of the frequency for which the determinant of the covariance matrix is greatest computed over the training space. The transformer inference module then attends over the inputs with the selected grid cell embeddings based on the DPP-A objective.
@@
The current approach has some limitations and presents interesting directions for future work. First, we derive the grid cell code from known properties of neural systems, rather than obtaining the code directly from real-world data. Here, we are encouraged by the body of work providing evidence for grid cell code in the hidden layers of neural networks in a variety of task contexts and architectures (Wei et al., 2015; Cueva & Wei, 2018; Banino et al., 2018; Whittington et al., 2020). This suggests reason for optimism that DPP-A may promote strong generalization in cases where grid cell code naturally emerge: for example, navigation tasks (Banino et al., 2018) and reasoning by transitive inference (Whittington et al., 2020). Integrating our approach with structured representations acquired from high-dimensional, naturalistic datasets remains a critical next step which would have significant potential for broader future practical applications. So too does application to more complex transformations beyond translation and scale, such as rotation, and complex forms of representations, and analogical reasoning tasks (Holyoak, 2012; Webb et al., 2023; Lu et al., 2022). Second, it is not clear how DPP-A might be implemented in a neural network. In that regard, Bozkurt et al. (2022) recently proposed a biologically plausible neural network algorithm using a weighted similarity matrix approach to implement a determinant maximization criterion, which is the core idea underlying the objective function we use for DPP-A (Equation 6), suggesting that the DPP-A mechanism we describe may also be biologically plausible. This could be tested experimentally by exposing individuals (e.g., rodents or humans) to a task that requires consistent exposure to a subregion, and evaluating the distribution of activity over the grid cells. Our model predicts that high frequency grid cells should increase their firing rate more than low frequency cells, since the high frequency grid cells maximize the determinant of the covariance matrix of the grid cell embeddings. It is also worth noting that Frankland & Cohen (2020) have suggested that the use of DPPs may also help explain a mutual exclusivity bias observed in human word learning and reasoning. While this is not direct evidence of biological plausibility, it is consistent with the idea that the human brain selects representations for processing that maximize the volume of the representational space, which can be achieved by maximizing the DPP-A objective function defined in Equation 6. Third, we compared grid cell code to only one-hots and place cell code. Future work could compare to a broader range of potential biological coding schemes for the overall space, e.g. boundary vector cell coding (Barry et al., 2006), band cell coding, or egocentric boundary cell coding (Hinman et al., 2019).
@@
Finally, we focus on analogies in linear spaces which limits the generality of our approach in nonlinear spaces. In that regard, there are at least two potential directions that could be pursued. One is to directly encode nonlinear structures (such as trees, rings, etc.) with grid cells, to which DPP-A could be applied as described in our model. The TEM model (Whittington et al., 2020) suggests that grid cells in the medial entorhinal may form a basis set that captures structural knowledge for such nonlinear spaces, such as social hierarchies and transitive inference when formalized as a connected graph. Another would be to use eigen-decomposition of the successor representation (Dayan, 1993), a learnable predictive representation of possible future states that has been shown by Stachenfeld et al. (2017) to provide an abstract structured representation of a space that is analogous to the grid cell code. This general-purpose mechanism could be applied to represent analogies in nonlinear spaces (Frankland et al., 2019), for which there may not be a clear factorization in terms of grid cells (i.e., distinct frequencies and multiple phases within each frequency). Since the DPP-A mechanism, as we have described it, requires representations to be factored in this way it would need to be modified for such purpose. Either of these approaches, if successful, would allow our model to be extended to domains containing nonlinear forms of structure. To the extent that different coding schemes (i.e., basis sets) are needed for different forms of structure, the question of how these are identified and engaged for use in a given setting is clearly an important one, that is not addressed by the current work. We imagine that this is likely subserved by monitoring and selection mechanisms proposed to underlie the capacity for selective attention and cognitive control (Shenhav et al., 2013), though the specific computational mechanisms that underlie this function remain an important direction for future research.